OpenStudy (anonymous):

Solve by factoring the trig equation 2sin^2-sinx-1=0

OpenStudy (cwrw238):

(2sinx + 1 )(sinx - 1 ) = 0 sin x = -1/2 or 1 x = 90, 210, 330 degrees ( in range 0 - 360)

OpenStudy (anonymous):

The answers need to be in terms of (0-2pi) I got 7pi/6, 2pi and 11pi/6, 2pi... Am i wrong?

OpenStudy (cwrw238):

90 degrees = pi/2, 210 degrees = 7pi/6, 330 degrees = 11pi/6 these are in radians

OpenStudy (anonymous):

Thank you! :)

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

2pi is incorrect